# The Past, Present, and Future of Automation in Model-Driven Engineering

### v1

@@ Abstract.
Model-Driven Engineering (MDE) provides a huge body of knowledge of automation for many different engineering tasks, especially those involving transitioning from design to implementation. With the huge progress made on Artificial Intelligence (AI) techniques, questions arise for the future of MDE such as how existing MDE techniques and technologies can be improved or how other activities which currently lack dedicated support can also be automated. However, at the same time, it has to be revisited where and how models should be used to keep the engineers in the loop for creating, operating, and maintaining complex systems. To trigger dedicated research on these open points, we discuss the history of automation in MDE and present perspectives on how automation in MDE can be further improved and which obstacles have to be overcome in the medium and long term perspective.
conference: International Workshop on Software Engineering in 2030; July 2024; doi:
@@ 1. Introduction
Bézivin coined the term Model-Driven Engineering (MDE) in 2005 as follows: “MDE is a software engineering methodology which uses formal models, i.e., models which are machine-readable and processable, to semi-automatically produce executable software systems” (Bézivin, 2005). The main idea behind formal models is to provide abstractions of software systems to manage their complexity. Based on these abstractions, as they are machine-readable and even processable, automation techniques are provided, e.g., to generate executable software (Schmidt, 2006), which was the main focus 20 years ago in the pioneering phase of MDE. However, soon, other automation concerns emerged and have been studied, such as analysing models by transforming them into formal domains such as model checkers. Model-to-model transformations have been used for this purpose. For code generation, model-to-code transformations have been developed to automate the transition from design models to code-based implementations. To complete the picture, code-to-model transformations have been used for reverse engineering of existing software systems such as legacy systems. A comprehensive study on different transformation intents is provided in (Lúcio et al., 2016).
@@
For developing automation support in terms of transformations, major research efforts have been spent on developing model transformation languages as it has been soon realized that transformations are complex software systems themselves (Sendall and Kozaczynski, 2003). In parallel, language engineering approaches based on metamodeling techniques have been emerging, which allow for the efficient and effective development of modeling languages and accompanying modeling tool support such as model editors, validators, and storage facilities (Atkinson and Kühne, 2003). Having these infrastructures enabled the development of domain-specific modeling languages (DSMLs) (Chen et al., 2005) in addition to general-purpose modeling languages such as UML.
@@
By having abstractions as models combined with automation support, the promise of MDE is to allow engineers to focus on the problem domain, i.e., which systems should be realized, rather than on the solution domain, i.e., how the system is realized. However, to realize this promise, several challenges have to be tackled as reported over the years in several papers. In (France and Rumpe, 2007), the authors discuss challenges associated with modeling languages, e.g., how to engineer DSMLs, achieving separation of concerns to deal with complexity, and model manipulation/management mostly supported by transformations. In (Mussbacher et al., 2014), the authors discuss how MDE may be used even beyond software, raising the need for interdisciplinary modeling support, e.g., for socio-technical systems. The latter requires model integration and consistency support as well as strong tool support. Finally, in a more recent work (Bucchiarone et al., 2020), the authors discuss important aspects of MDE such as foundational, social, community, domain, and tool aspects. However, we are not aware of any report focusing explicitly on the automation aspect of MDE and how this aspect will develop towards 2030. It is important to close this gap as with the recent and expected advances of AI, the current automation techniques and technologies of MDE are highly impacted as well and new ones may be developed to support additional tasks in the software engineering process which are not targeted by existing technologies. Thus, a discussion on the role of models, MDE methodologies, and engineers in this new era is required to frame the potential future of MDE.
@@
Important questions have to be answered, such as which phases can be additionally automated, e.g., should we directly go from natural language requirements to code, or is there still, especially because of the enhanced AI capabilities, a sweet spot for using models in such highly automated processes? Can we now ease and automate the modeling process itself and reach at the same time a higher automation degree in the synthesis process, such as code generation? These questions are important for the foundations of MDE, but at the same time, also for recent application areas such as low-code/no-code development (Ruscio et al., 2022) and digital twin engineering (Tao et al., 2018).
In the following, we capture the achievements of automation in MDE by enumerating prominent engineering activities which have been in the focus of automation as well as the automation techniques which have been applied for achieving the automation. Based on this foundation, we lay out several challenges hindering further significant automation increases and conclude the paper with perspectives on tackling these challenges until 2030.
@@ 2. Achievements in Automation
2.1. Targeted Automation Activities
In this section, we categorize and discuss the activities that have been targeted for automation within the MDE field, organizing them according to the categories outlined in the contents for a Model-Based Software Engineering Body of Knowledge (Burgueño et al., 2019).
@@ 2.1.1. Model Quality Management
Significant efforts have been made to automate activities for finding defects, faults and failures in models and modeling processes. Numerous model validation and verification techniques exist that automate quality checks, such as correct syntax, semantics, structure, behaviour, etc., of models to ensure that any inconsistencies or errors are detected and rectified at the model level and as soon as possible. For instance, model checking has been automated through algorithms that check the temporal behaviour of a software or system model against certain properties or specifications, ensuring they meet specified requirements or constraints, and there are numerous languages and mature tools available (Holzmann, 2004; Bengtsson et al., 1995; Kwiatkowska et al., 2011).
Automated testing techniques based on models (a.k.a. model-based testing) have been developed, too, where models are used to generate test cases automatically (e.g., (Pretschner et al., 2005)). Traditionally, this method has helped identify potential problems early in the development process, significantly reducing cost and effort in the testing phase. Moreover, in recent years, model-based testing automation has gained relevance due to the rise of digital twins (Bickford et al., 2020), which impose the need for streamlining testing, and the adoption of new development methodologies such as agile, continuous development, integration and delivery (Jongeling et al., 2019) where the process of executing test cases and analyzing the results, often in a continuous integration / continuous deployment (CI/CD) pipeline that needs to be automated.
@@ 2.1.2. Model-based execution
For testing, debugging or analysis, executable models and simulation facilities are necessary. Regarding executable models (Kleppe et al., 2003), executable UML models are popular, which usually use an action language to define the behaviour of the system (e.g., the Alf Action Language (alf, 2017), SOIL (Schäfer and Gogolla, 2020), or FUML [^1] https://www.omg.org/spec/FUML).
Over the years, the automation of model-based execution, simulation, and co-simulation has seen numerous advancements, moving from manual processes that required human intervention—e.g., to set up simulations and adjust parameters based on expert knowledge and observations—to highly automated processes and tools. For instance, there has been the standardization of an Action Language as part of the OMG UML Specification for the execution of UML models (alf, 2017) and tools for simulation such as Modelica [^2] https://modelica.org, which are widely used by system engineers. There is also available a series of interpreters for modeling languages, both graphical and textual, many of them fully integrated into modeling environments, e.g., see (Gogolla et al., 2007).
[…]
@@ 2.1.3. Model Transformations
Model transformations constitute a fundamental aspect of MDE and serve as the backbone for automating the transition from high-level abstractions to executable code, simulations, or other models. They can be classified into Model-to-model transformations (M2M), Model-to-text (M2T) transformations, and Text-to-model (T2M) transformations (Kleppe et al., 2003). M2M transformations are perhaps the most exploited automation and automated activity in MDE, where, among other uses, models are automatically transformed into executable code in various programming languages. M2T transformations provide the ability to autogenerate code from models, contributing significantly to software development automation. The models serve as the primary artifact; from them, code, documentation or other textual artifacts (such as JSON files) are generated automatically (Klein, 2015). T2M transformations are used for reverse-engineering code into models as well as other purposes such as documentation generation (Wang et al., 2010).
@@
At the same time that model transformations have enabled and pushed forward automation in software engineering, they have also been the target of automation. This automation leverages specific languages, execution engines, tools, and frameworks, as described below. One of the primary means of automating model transformations is through dedicated transformation languages. These languages provide constructs tailored to describe how model elements must be mapped to their equivalent elements. Examples of such languages include ATL (Jouault et al., 2008), QVT (OMG, 2005), ETL (Kolovos et al., 2008), and Henshin (Arendt et al., 2010) to name just a few—for a survey see (Kahani et al., 2019).
@@
Transformation engines [^4] https://projects.eclipse.org/projects/modeling.mmt.qvt-oml [^5] https://wiki.eclipse.org/ATL/EMFTVM(Burgueño et al., 2016; Cuadrado et al., 2022; Tisi et al., 2013) are specialized software components that execute the transformation rules defined in a transformation language against source models to generate target models. These engines manage the complexities of navigating model structures, applying transformation rules, resolving dependencies, and ensuring the integrity of the generated models. Transformation engines can be standalone tools or integrated into broader MDE platforms, providing a seamless experience from model editing through to transformation.
@@
For automation to be truly effective, model transformations must be well-integrated with the software development environment. This includes providing user-friendly interfaces for defining and managing transformations. Such integration enables developers to incorporate model transformations seamlessly into their workflows, automating repetitive tasks, and facilitating more agile and iterative development processes. Some efforts include the Epsilon Playground (Kolovos and García-Domínguez, 2022) and AToMPM (Syriani et al., 2013) among others.
@@ 2.1.4. Model Management
Model management involves different activities, which we discuss in the following. Automated model evolution involves strategies for evolving models to reflect changes in requirements or design. Automation can be achieved through predefined transformation rules or templates that apply changes across models, ensuring consistency and reducing manual effort. Languages such as those that conform to Epsilon (Kolovos and García-Domínguez, 2022) or ATL (Jouault et al., 2008) support automated transformations that facilitate model evolution.
When metamodels evolve, models that conform to those metamodels may become invalid and need to be adapted to remain consistent with the new version of the metamodel (Herrmannsdoerfer and Ratiu, 2009). Automated model migration is needed to update models. So far, this has been achieved in numerous ways, for instance, by means of model transformations (Sprinkle and Karsai, 2004; Ruscio et al., 2012) or with dedicated languages such as Epsilon Flock (Rose et al., 2014).
@@
Since models are software artifacts, version control systems are beneficial to keep up with current development needs. Given the different nature of models w.r.t. code, approaches to deal with models are needed. For instance, EMFStore (Koegel and Helming, 2010) or Connected Data Objects (CDO) [^6] https://www.eclipse.org/cdo provide automated support for model versioning.
@@
Refactoring models is the process of restructuring the model while not modifying its original functionality or behaviour. This process can be automated using model transformation languages and refactoring libraries. Tools and frameworks, such as Epsilon (Kolovos and García-Domínguez, 2022) or EMF Refactor (Arendt and Taentzer, 2012),provide automated support for model refactoring.
2.1.5. Modeling
Last but not least, the task of modeling itself, i.e., creating models as abstractions of software and systems, has also been the target of automation. In particular, specific efforts have been put into approaches for model generation (Saini et al., 2022), model search (López and Cuadrado, 2020), model completion (Chaaben et al., 2023) and model classification (Nguyen et al., 2019). Among other efforts, a long-lasting line of work to automate the execution of models has been heavily addressed by researchers working on models@runtime (Bencomo et al., 2019), where the models of a system need to be up-to-date while the system is running at the same time that is used to adapt the runtime behaviour.

### v2

# Automation in Model-Driven Engineering: A look back, and ahead

@@ Abstract.
Model-Driven Engineering (MDE) provides a huge body of knowledge of automation for many different engineering tasks, especially those involving transitioning from design to implementation. With the huge progress made in Artificial Intelligence (AI), questions arise about the future of MDE, such as how existing MDE techniques and technologies can be improved or how other activities that currently lack dedicated support can also be automated. However, at the same time, it has to be revisited where and how models should be used to keep the engineers in the loop for creating, operating, and maintaining complex systems. To trigger dedicated research on these open points, we discuss the history of automation in MDE and present perspectives on how automation in MDE can be further improved and which obstacles have to be overcome in both the medium and long-term.
conference: International Workshop on Software Engineering in 2030; November 2024; Puerto Galinàs (Brazil)copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Bézivin defined the term Model-Driven Engineering (MDE) in 2005 as follows: “MDE is a software engineering methodology which uses formal models, i.e., models which are machine-readable and processable, to produce executable software systems semi-automatically” (Bézivin, 2005). The main idea behind formal models is to provide abstractions of software systems to manage their complexity. Based on these abstractions, as they are machine-readable and even processable, automation techniques are provided, e.g., to generate executable software (Schmidt, 2006), which was the main focus 20 years ago in the pioneering phases of MDE. However, other automation concerns soon emerged and have been studied, such as analysing models by transforming them into formal domains such as model checkers. Model-to-model transformations have been used for this purpose. Model-to-code transformations have been developed to automate the transition from design models to code-based implementations. To complete the picture, code-to-model transformations have been used to reverse engineer existing software systems such as legacy systems. A comprehensive study on different transformation intents is provided in (Lúcio et al., 2016).
@@
Soon the community realized that transformations are complex software systems themselves (Sendall and Kozaczynski, 2003). Hence, significant research efforts have been spent developing model transformation languages to develop automation support in terms of transformations. In parallel, language engineering approaches based on metamodeling techniques (Atkinson and Kühne, 2003) have been emerging, which allow for the efficient and effective development of modeling languages and accompanying modeling tool support such as model editors, validators, and storage facilities. These infrastructures enabled the development of domain-specific modeling languages (DSMLs) (Chen et al., 2005) in addition to general-purpose modeling languages such as UML.
@@
By having abstractions as models combined with automation support, the promise of MDE is to allow engineers to focus on the problem domain, i.e., which systems should be realized, rather than on the solution domain, i.e., how the system is realized. Achieving this promise involves addressing several challenges, which have been discussed in various studies. In (France and Rumpe, 2007), the authors discuss challenges associated with modeling languages, e.g., how to engineer DSMLs, achieving separation of concerns to deal with complexity, and model manipulation/management mostly supported by transformations. In (Mussbacher et al., 2014), the authors discuss how MDE may be used even beyond software, raising the need for interdisciplinary modeling support, e.g., for socio-technical systems. The latter requires model integration, consistency, and strong tool support. Finally, in a more recent work (Bucchiarone et al., 2020), the authors discuss important aspects of MDE, such as foundational, social, community, domain, and tool aspects. However, we are not aware of any report explicitly focusing on the automation aspect of MDE and how this aspect will develop toward 2030. It is important to close this gap as with the recent and expected advances of AI, the current automation techniques and technologies of MDE are highly impacted as well, and new ones may be developed to support additional tasks in the software engineering process that are not targeted by existing technologies. Thus, a discussion on the role of models, MDE methodologies, and engineers in this new era is required to frame the potential future of MDE. This is also of paramount importance for the further adoption of MDE in different domains. Previous studies have reported successful adoptions of MDE for particular domains (e.g., see (Vallecillo, 2015; Liebel et al., 2018; Nistala et al., 2024; Daniel and Cabot, 2024; Buschhaus et al., 2024; Sadovykh et al., 2024; James et al., 2024; Kurtev et al., 2024)). However, challenges and needs for the adoption of MDE have been reported (Hutchinson et al., 2014; Holtmann et al., 2024) recently, which may be tackled with improved automation capabilities such as improving the interoperability of MDE tools as well as providing sophisticated model management support to allow for agile processes. Moreover, for successful adoption, it is essential to rely on high-quality models (in terms of completeness, correctness, and consistency with the systems realized) that evolve to reach high automation levels (Holtmann et al., 2024). Of course, reaching these high-quality levels requires additional efforts, which may be mitigated by the application of AI in the future.
@@
To tackle these challenges, important questions have to be answered, such as which phases can be additionally automated, e.g., should we directly go from natural language requirements to code, or is there still, mainly because of the enhanced AI capabilities, a sweet spot for using models in such highly automated processes? Can we now ease and automate the modeling process itself and reach, at the same time, a higher degree of automation in the synthesis process, such as code generation? These questions are important for the foundations of MDE, but at the same time, also for recent application areas such as low-code/no-code development (Ruscio et al., 2022) and digital twin engineering (Tao et al., 2018).
This article pursues two primary objectives. The first is to conduct a retrospective analysis of how automation needs in Model-Driven Engineering (MDE) have historically been addressed through the application of various enabling technologies. The second objective is to explore future perspectives on MDE automation. These forward-looking considerations are influenced by two key factors. First, the evolution and transformation of existing paradigms, such as the rapid advancements in machine learning and artificial intelligence, are redefining the potential for automation. Second, the emergence of novel paradigms, like quantum computing or neuromorphic computing, presents unprecedented opportunities to rethink the scope and scale of automation in MDE. Additionally, the growing emphasis on non-functional concerns, including security, sustainability, and ethical considerations, introduces new dimensions to how automation solutions are envisioned and implemented. It is important to highlight that future perspectives on automation are not merely responses to the limitations of previous approaches. Instead, they are driven by the creation of new opportunities arising from the continuous evolution of technology and the emergence of cutting-edge paradigms. This dual focus on past achievements and future possibilities aims to inspire innovative directions for advancing automation in MDE.
To achieve these objectives, we began by surveying key activities in MDE that require automation (Section 2.1). Subsequently, we identified the primary technologies that have enabled varying degrees of automation for these activities and gathered illustrative examples of representative work (Section 2.2). While we collected a substantial body of work, it is important to note that we did not conduct a systematic literature review, as this would extend beyond the scope of our article. The third step involved synthesizing our findings by mapping tasks to automation technologies, thereby completing the overall picture (Section 2.3). For the exploration of future automation perspectives until 2030, we examined both challenges and opportunities across four critical dimensions: advancements in artificial intelligence, the role of human factors, the diversity of application domains, users, and modeling interfaces, and the growing importance of non-functional concerns such as security and sustainability (Sections 3.1 to 3.4).
@@ 2. Achievements in Automation
In this section, we present the evolution of MDE automation over the years. We divide it into two subsections: Section 2.1 presents the activities and MDE that require automation, and Section 2.2 presents the primary technologies that have enabled such automation together with examples of them..
2.1. Targeted Automation Activities [^1] Note that we will refer to activities as activities or as tasks indifferently.
In this section, we categorize and discuss the activities that have been targeted for automation within the MDE field, organizing them according to the categories outlined in the contents of a Model-Based Software Engineering Body of Knowledge (Burgueño et al., 2019). It should be noted that while these basic tasks serve as foundational components, they can be composed into more complex automation pipelines for different purposes.
2.1.1. Modeling
The task of modeling itself, i.e., creating models as abstractions of software and systems, has been the target of automation for many years. In particular, specific efforts have been put into approaches for the creation of DSLs and DSVLs, for instance, using frameworks and tools such as JetBrains MPS [^3] https://www.jetbrains.com/mps or the Eclipse Modeling Framework (EMF) [^4] https://eclipse.dev/modeling/emf. Modeling tools have also helped improve MDE workflows and reduce the manual effort required. Examples of these tools are Papyrus [^5] https://eclipse.dev/papyrus and MagicDraw [^6] https://www.magicdraw.com .
Other modeling activities that over the years have been the target of automation are model generation (Saini et al., 2022), model search (López and Cuadrado, 2020), model completion (Chaaben et al., 2023) and model classification (Nguyen et al., 2019). A long-lasting line of work to automate the execution of models, which is worth emphasizing, has been heavily addressed by researchers working on models@runtime (Bencomo et al., 2019), where the models of a system need to be up-to-date while the system is running at the same time that is used to adapt the runtime behavior. This line of work is a prominent example of having automated modeling in use.
@@ 2.1.2. Model Quality Management
Significant efforts have been made to automate activities to find defects, faults, and failures in models and modeling processes. Numerous model validation and verification techniques exist that automate quality checks, such as correct syntax, semantics, structure, behavior, etc., of models to ensure that any inconsistencies or errors are detected and rectified at the model level and as soon as possible. For instance, model checking has been automated through algorithms that check the temporal behavior of a software or system model against certain properties or specifications, ensuring they meet specified requirements or constraints, and there are numerous languages and mature tools available (Holzmann, 2004; Bengtsson et al., 1995; Kwiatkowska et al., 2011).
Automated testing techniques based on models (a.k.a. model-based testing) have been developed, too, where models are used to generate test cases automatically (e.g., (Pretschner et al., 2005)). Traditionally, this method has helped identify potential problems early in the development process, significantly reducing cost and effort in the testing phase. Moreover, in recent years, model-based testing automation has gained relevance due to the rise of digital twins (Bickford et al., 2020), which impose the need for streamlining testing, and the adoption of new development methodologies such as agile, continuous development, integration, and delivery (Jongeling et al., 2019) where the process of executing test cases and analyzing the results needs to be automated in continuous integration/continuous deployment (CI/CD) pipelines.
@@ 2.1.3. Model-based execution
It is necessary to have executable models (Kleppe et al., 2003) and simulation facilities for testing, debugging, or analysis purposes. In this regard, executable UML models are popular and usually use action languages to define the behavior of the system (e.g., SOIL (Schäfer and Gogolla, 2020), or FUML [^7] https://www.omg.org/spec/FUML).
Over the years, the automation of model-based execution, simulation, and co-simulation have seen numerous advancements, moving from manual processes that required human intervention—e.g., to set up simulations and adjust parameters based on expert knowledge and observations—to highly automated processes and tools. For instance, there has been the standardization of an Action Language as part of the OMG UML Specification for the execution of UML models (OMG, 2017) and tools for simulation such as Modelica [^8] https://modelica.org, which are widely used by system engineers. There is also available a series of interpreters for modeling languages, both graphical and textual, many of them fully integrated into modeling environments, e.g., as in the case of the USE approach (Gogolla et al., 2007).
Furthermore, there have been efforts to come up with model execution support by developing dedicated interpreters for domain-specific modeling languages, e.g., see initiatives such as GeMOC (Globalization of Modeling Languages) [^9] https://gemoc.org focusing on the development of technologies and methodologies to generate dedicated execution-oriented tools for models such as model testing and debugging capabilities based on the defined operational semantics of the modeling languages.
@@ 2.1.4. Model Transformations
Model transformations constitute a fundamental aspect of MDE and serve as the backbone for automating the transition from high-level abstractions to executable code, simulations, or other models. These are perhaps the most exploited automation and automated activity in MDE, where, among other uses, models are automatically transformed into executable code in various programming languages. They can be classified into model-to-model transformations (M2M), model-to-text (M2T) transformations, and text-to-model (T2M) transformations (Kleppe et al., 2003). M2M transformations receive a model as input and provide another model as output (e.g., a BPMN model can be transformed into a Petri Net model). M2T transformations provide the ability to autogenerate text from models, contributing significantly to software development automation, where the models serve as the primary artifact and, from them, textual artifacts are automatically generated (Klein, 2015) such as code (e.g., C# code from UML state machines), documentation, or other artifacts (such as JSON files for data exchange). T2M transformations are used to reverse engineer code in models, as well as other purposes such as documentation generation (Wang et al., 2010) (for example, to convert legacy code into a software architecture model).
@@
At the same time, model transformations have enabled and pushed forward automation in software engineering; they have also been the target of automation. This automation leverages specific languages, execution engines, tools, and frameworks, as described below. One of the primary means of automating model transformations is through dedicated transformation languages. These languages provide constructs that describe how model elements must be mapped to their equivalent elements. Examples of such languages include ATL (Jouault et al., 2008), QVT (OMG, 2005), ETL (Kolovos et al., 2008), and Henshin (Arendt et al., 2010) to name just a few—for a survey see (Kahani et al., 2019).
@@
Transformation engines [^10] https://projects.eclipse.org/projects/modeling.mmt.qvt-oml, [^11] https://wiki.eclipse.org/ATL/EMFTVM(Burgueño et al., 2016; Cuadrado et al., 2022; Tisi et al., 2013) are specialized software components that execute the transformation rules defined in a transformation language against source models to generate target ones. These engines manage the complexities of navigating model structures, applying transformation rules, resolving dependencies, and ensuring the integrity of the generated models. Transformation engines can be standalone tools or integrated into broader MDE platforms, providing a seamless experience from model editing to transformation.
@@
Model transformations must be well-integrated with the software development environment for automation to be truly effective. This includes providing user-friendly interfaces for defining and managing transformations. Such integration enables developers to incorporate model transformations seamlessly into their workflows, automating repetitive tasks and facilitating more agile and iterative development processes. Some efforts include the Epsilon Playground (Kolovos and García-Domínguez, 2022) and AToMPM (Syriani et al., 2013), among others.
@@ 2.1.5. Model Management
Model management involves different activities, which we discuss in the following. Automated model evolution involves strategies for evolving models to reflect changes in requirements or design. Automation can be achieved through predefined transformation rules or templates that apply changes across models, ensuring consistency and reducing manual effort. Languages such as Epsilon (Kolovos and García-Domínguez, 2022) or ATL (Jouault et al., 2008) support automated transformations that facilitate model evolution.
When metamodels evolve, models that conform to those metamodels may become invalid and must be adapted to remain consistent with the new version of the metamodel (Herrmannsdoerfer and Ratiu, 2009). Automated model migration is needed to update models. So far, this has been achieved in numerous ways, for instance, by means of model transformations (Sprinkle and Karsai, 2004; Ruscio et al., 2012) or with dedicated languages such as Epsilon Flock (Rose et al., 2014).
@@
Since models are software artifacts, version control systems are beneficial for keeping up with current development needs. Given the different nature of models w.r.t. code, approaches to deal with models are needed. For instance, EMFStore (Koegel and Helming, 2010) or Connected Data Objects (CDO) [^12] https://www.eclipse.org/cdo provide automated support for model versioning.
@@
Refactoring models is the process of restructuring the model while not modifying its original functionality or behavior. This process can be automated using model transformation languages and refactoring libraries. Tools and frameworks, such as Epsilon (Kolovos and García-Domínguez, 2022) or EMF Refactor (Arendt and Taentzer, 2012), provide automated support for model refactoring.